OpenStudy (cedrick12345):

Find the complete factored form of the polynomial: w^6x^2y^4-w^4xy

OpenStudy (michele_laino):

hint: we can factor out \(w^4xy\)

OpenStudy (michele_laino):

namely, we can write this: \[\huge {w^4}xy\left( {term\;1 - term\;2} \right)\]

OpenStudy (michele_laino):

hint: we have this: \[\huge \begin{gathered} term\;1 = \frac{{{w^6}{x^2}{y^4}}}{{{w^4}xy}} = ... \hfill \\ \hfill \\ term\;2 = \frac{{{w^4}xy}}{{{w^4}xy}} = ... \hfill \\ \end{gathered} \]

OpenStudy (cedrick12345):

what